[figure: 9.7 cm]

PHASE FLUCTUATIONS IN SHORT COHERENCE LENGTH SUPERCONDUCTORS

B. Tobijaszewska and R. Micnas Institute of Physics, A. Mickiewicz University, Umultowska 85, 61-614 Poznan, Poland

We analyze the effects of phase fluctuations on the extended s-wave and d-wave superconductivity in 2D model with intersite electron pairing described by the extended Hubbard Hamiltonian with the nearest–neighbour and next–nearest–neighbour hopping. Within the Kosterlitz-Thouless scenario the critical temperatures were determined and compared with the ones from BCS-HFA. Our results indicate a possibility of new phases in which pairs are incoherent, but the fermionic spectrum has a gap. The Uemura-type plots, i.e. the critical temperature vs zero temperature phase stiffness, were obtained for extended s-wave and d-wave pairings.

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : POSTER

Corresponding author : B. Tobijaszewska

Address for correspondence : Institute of Physics, A. Mickiewicz University, Umultowska 85, 61-614 Poznan, Poland

Email address : bch@delta.amu.edu.pl